OpenStudy (anonymous):

John is now 3 times as old as his brother Sam. In 5 years, John will be twice as old as Sam will be then. Find their present ages John is now 3 times as old as his brother Sam. In 5 years, John will be twice as old as Sam will be then. Find their present ages @Mathematics

OpenStudy (anonymous):

Sam is x John is 3x In 5 years Sam is x + 5 John is 3x + 5

OpenStudy (anonymous):

So the equation is 3x + 5 = 2(x + 5)

OpenStudy (anonymous):

3x + 5 = 2x + 10 x = 5

OpenStudy (anonymous):

So you go back to Sam = x John = 3x and plug in x = 5 Sam = 5 John = 15

OpenStudy (anonymous):

THANKS GUY